OpenStudy (anonymous):

Solve for x. 8(x + 1) - 3(x + 4) = 7(2 - x)

OpenStudy (anonymous):

8(x+1) = 8x + 8 -3(x + 4) = -3x - 12 7(2 - x) = 14 - 7x 8x + 8 - 3x - 12 = 14 - 7x 5x - 4 = 14 - 7x Add 7x to both sides 8x + 7x - 4 = 14 - 7x + 7x 15x - 4 = 14 add 4 to both sides 15x - 4 + 4 = 14 + 4 15x = 18 divide both sides by 15 15x/15 = 18/15 x = 18/15 = 6*3/3*5 = 6/5 Source(s): Longtime college math teacher 4 years ago

OpenStudy (anonymous):

All of the choices are mixed fractions

OpenStudy (anonymous):